Minutes of the Tracker Meeting

October 7, 2004


See the lists of actions below.


(1) Progress on Tower-0 & plan for debugging the bad cable

Emilio reported that the Tower-0 assembly and alignment activities are finished.  A list of open issues was sent to SLAC.  One issue was that it was impossible to install any of the 8 studs for the side flexures because of the Nytemp patches that provide such a good locking feature that it is impossible to get the nuts threaded on either end.

Today they started running the comprehensive performance test (CPT) in the tower upright position.  The top cable bending was done on Tuesday, and the good news is that it did not damage anything.

Luca reported on the CPT, which was started today.  So far they measured the leakage current on only one side: 20uA without the dark box.  They saw no change in current with the lights off.  The current matches expectations.  They will repeat this test on all 4 sides. One goal of the present CPT is to check the execution time, in addition to the performance tests.  Luca said that the really good news is that there was no change in the incidence of dead or noisy channels during the assembly.  The trays all survived the handling without incident.

There is still the problem that cable 3 doesn’t function.  In general they get no data back from even the Layer-0 GTRC, although there was a mention that occasionally the GTFE register load and read works.  They are able to write to the Layer-0 GTRC, because they can write registers via that GTRC and then read them from the opposite side.  They will extend this test to all layers.  Robert suggested that the problem must lie with the data line between the Layer-0 GTRC on Cable 3 and the TEM.  It could also be a problem with the data-line bias on the top-most GTRC.  However, the flight cables bias that input via resistors in addition to and redundant to the internal self-bias of the GTRC data receiver, so it seems unlikely that both would fail.  Robert suggested installing a breakout board and looking with a scope at the data coming back from the GTRCs.

Robert requested the Pisa find a way to test the application of Solithane to sidewall fasteners before Tower-A assembly starts.  No Solithane was used on Tower-0.


(2) Tower A preparations & issues

We went through the Power-Point file of issues that Emilio sent.  The following is some of the points discussed.

(a) Problems with INFN parts and components

-       Inspect the bare trays for glue in the insert threads both at Plyform and at Pisa during ESPI testing, so that they can be cleaned out before G&A integrates the electronics and SSDs.

(b) Problems with SLAC parts and components

-       Implement a thorough 100% testing of the studs+Nytemp before delivery to Pisa

-       The flight-article tie-down straps are coming in very soon and will be CMM’ed and reviewed to print.

-       SLAC should change the torque spec for the connector mount to 0.4N-m.  The present spec is dangerous.

-       Pisa should send a redline drawing showing where to add a grounding screw to the tower baseplate.

-       SLAC should check whether nest fasteners can be obtained with a locking feature suitable for thermal/vacuum testing.  Also, the correct fastener with the button head should be sent.

(c) Electrical test issues

-       Pisa will modify some Micro-D connector savers so that they stay screwed into the TEM.

(d) Cable assembly and bending

-       The cable holding plate design needs to be reviewed with Martin.  It appears that the cable offsets don’t work out such that the cables line up with the connection points without twisting. 

-       INFN wants to put some foam in place during the vibration testing to protect the cables.  Jeff will look for some suitable foam.

-       The tape for the cables really doesn’t work.  It is too thin and peels off very easily, and it does not fill the gap between tray and sidewall.  Pisa is waiting on a bid for some O-ring material in Viton or Silicone, to try to implement Sandro’s idea.  Jeff will also look for suitable material.  Robert needs to talk to Martin about this design problem.  Emilio will send the dimensions needed for the O-ring material.

-       Arthur needs to correct the bending tools, make more copies, and make a special tool for the uppermost cactus arm.

-       Pisa would like the screw hole in the cable clip to be smaller, so that the screw can be kept away from the cable edges during installation.  Also, they do not want to put tape on the bracket.  It does not help hold the cable down, and it makes it hard to adjust the cable position side to side.

-       Steve will check on the fit of the –x,-y corner bracket.

-       Jeff will try to get a long (2-inch) screwdriver bit for the clip installation.

-       Jeff will get the clip’s edges rounded over and will get Kapton tape to coat the inner face of the clip.


(3) MCMs to return to SLAC for inspection/rework

Luca reported that there still is 1 MCM at G&A without 75-ohm clock-bus termination resistors.  It will be returned to SLAC for rework.  Also, Robert requested that all unused flight MCMs that went to Pisa before the latest shipment of 40 be returned to SLAC, along with the one MCM in the 40 that was accidentally shipped with an open NCR.  The reason is that the latest shipment went through additional inspection and screening steps that were not applied to previous MCMs.  These include the conformal coating inspections, inspection of pitch adapters for broken traces after the environmental testing, and screening out of MCMs with low margin with respect to clock duty cycle.


Action Items from Last Week:

- Arthur: write and send out a summary of the baseplate deformation issue.

- Arthur: contact QA people at Nyloc.

- Jeff: get the remaining studs dycronited asap.

- Jeff: procure a better socket wrench for stud installation. It needs to be longer with a tighter fit.

- Peter: get more extraction tools to Pisa for removing cones with the base plate in place.

- Dave: set up a meeting on the baseplate deformation issue.

- Sandro: update the 1584 procedure with references to the new procedures.

- Dave: set up a call-in for a G&A MRB at 9:00 am Wednesday.

- Luca: hand carry MCMs to Pisa Saturday (flight and non-flight).

- Dave: get a high-level meeting going to push Parlex to perform.

- Jeff: get 10 spare square nuts for test work at Nytemp.

- Jeff: expedite an order for new cable clips for the cable holding fixture.